OpenStudy (anonymous):

A diamond is purchased for $3500 It's value increases %5 each year. Find the value of the diamond after 6 years.

OpenStudy (anonymous):

\[3500(1.05)^6\] and a calculator

OpenStudy (anonymous):

to increase a number by 5% multiply it by \(1.05\) to do that six times, multiply by \((1.05)^6\)

OpenStudy (jiteshmeghwal9):

The value of diamond is $3500. so, 5% of $3500=5/100*3500=$175 now it increases $175 each year. so, after 6 years it will be (6*175)+3500=?